# TTSnap: Test-Time Scaling of Diffusion Models via Noise-Aware Pruning

### v1

@@ 1 Introduction
Figure 1: Comparison of the best-of-NN TTS algorithm with our TTSnap algorithm under an equivalent compute budget. By applying early-stage pruning, TTSnap can explore a much larger pool of candidates (shown as polygons). Candidate colors indicate reward values from low (light yellow) to high (dark red). Each candidate is denoised from left to right under different initial noise. The magnifying glass represents the reward models, where the noise-aware variants are finetuned using our self-distillation strategy.
@@ 3 Preliminaries
and may not be differentiable with respect to the inference trajectory. The search algorithm aims to identify the most promising candidates that maximize the verifier’s reward scores. The algorithm $f$ takes the verifier $\psi\in\Psi$, the diffusion model $\phi\in\Phi$, and NN candidate image samples with text conditions as input and returns an image:
@@
Increasing the number of candidates NN increases the probability of finding a high-reward sample, but also increases the inference-time compute cost. Among various search algorithms, best-of-NN is the most straightforward yet effective one [sana1_5, tts_ma]. This algorithm independently samples NN Gaussian noise, denoises each, and verifies their rewards. Its core advantage lies in enabling global exploration across the initial noise space. Other strategies, such as greedy neighborhood search [tts_ma, tts_greedy] or trajectory-level search [fk, tts_ma, dynamic_search, beam_vseq, tts_kim], primarily focus on finding locally optimal candidates. In this paper, we mainly focus on improving TTS over best-of-NN, since a more global search, as discussed in Sec. 6.3, typically yields better performance. While some methods scale performance by refining the conditioning signals [reflect_perfect, reflectdit, tts_kim_p], we focus our study to the setting where the text condition remains fixed.
@@ 4 The TTSnap Algorithm
The primary limitation of the best-of-NN strategy lies in its high verification cost, as each candidate must be fully denoised before reward evaluation. This constraint severely restricts the breadth of candidate exploration under a fixed computational budget. To address this inefficiency, we propose a pruning strategy, outlined in Algorithm 1, that verifies and eliminates unpromising candidates early in the denoising process. Since the reward models operate in the image domain, we approximate the image corresponding to an intermediate latent by estimating its posterior mean, detailed in Sec. 3, and decoding it via the variational auto-encoder (VAE) decoder. The intermediate images are then scored by the reward model, and candidates with low rewards are pruned according to a predefined retention ratio. This process is iterated until the final step, where the remaining candidates are fully denoised, and the one yielding the highest reward is selected as the output. The set of verification timesteps is denoted as $\mathcal{T}=\{\tau_{1},\dots,\tau_{m-1}\}\subseteq[1,M-1]$, where $\tau_{m-1}=M-1$. $M$ is the final timestep. The set of corresponding retention ratios is $\mathcal{A}=\{\alpha_{1},\dots,\alpha_{m-1}\}$, where $0<\alpha_{i}<1$ is the fraction of candidates that will be retained at timestep $\tau_{i}$.
@@ 12    $\mathcal{X}\leftarrow\{\mathcal{X}_{i}\mid i\in\operatorname*{arg\,topk}(\mathcal{R},k)\}$
By avoiding the full denoising of every candidate, such an algorithm can substantially reduce the overall compute cost. However, it introduces additional verification overhead: VAE decoding and reward computation. Let the computational cost of a single denoising step for one candidate be $B_{d}$, and that of a single verification be $B_{v}$. For NN initial candidates, the total compute budget $B$ can be expressed as
@@
Under the same compute budget $B$, our NN is much larger than that of the original best-of-NN strategy: $N=\lfloor B/(MB_{d}+B_{v})\rfloor$.
@@
The primary goal of TTS is to maximize the generated image quality with minimal computational cost. Existing TTS methods are typically evaluated under a fixed computation budget or by visualizing the relationship between the final reward and computational cost [dynamic_search, tts_ma, reflectdit]. Such evaluations are often indirect. Hence, we introduce a new metric that quantifies performance improvement as the computation budget increases. Let $f(N)$ denote the final image obtained from algorithm $f$ with NN noise candidates, as described in Eq. 6, and let $\psi(f(N))$ represent its corresponding reward score. The number of explored candidates NN can be determined from a given computation budget $B$, such as Eq. 8 for our method. Over a budget range $B^{-},B^{+}]$, we define the incremental performance gain with respect to budget $B$ as
@@ Results.
We construct a validation set using 200 prompts unseen during training, generating 80 samples per prompt along with all corresponding intermediate images. For the purpose of TTS, our goal is to ensure that reward predictions at early inference stages preserve a ranking similar to the final clean-image counterparts. To measure this rank consistency, we compute Kendall’s $\tau$ coefficient [kendall] between the noise-aware reward scores at early timesteps and the final-step reward scores for each prompt, and report the average correlation across all prompts on the validation set. Kendall’s $\tau$ measures the ordinal agreement between two ranked lists, taking values from 11 (identical order) to 1-1 (complete reversal). It reflects the normalized number of pairwise swaps needed to turn one ranking into the other, independent of the value magnitudes. As shown in Fig. 3, our noise-aware reward self-distillation consistently improves early-stage ranking agreement across all three reward models compared with their un-finetuned baselines. This substantial gain in rank consistency makes our early-step pruning framework more reliable and effective. A further analysis of NARF and a comparison with its alternatives are presented in Appendix Appendix B.
@@ 6.2 TTSnap Experiments
We take the fine-tuned noise-aware reward model from the previous section and evaluate our TTSnap algorithm on the validation prompts. The computation budget is measured in TFLOPs. We cap the maximum budget at 4,000 TFLOPs, which approximately equals 400 forward passes of the Flux diffusion transformer. Our main evaluation metric is $w$ as defined in Eq. 11, representing the increase ratio of the integral of reward improvement over budget, with reference to the best-of-NN baseline. To mitigate randomness and ensure a fair comparison across different TTSnap configurations, we generate a pool of over 200 candidate trajectories for each prompt and constrain all search procedures to operate within this shared pool. For each search, we iterate 100 times over different seeds and report the average reward of the final selected image.
@@
Figure 4: Analysis of the effect of different variables on the PickScore reward or number of explored samples NN, with respect to (a) NN, (b) the compute budget $B$, (c) the retention ratio $\alpha$, and (d) the pruning timestep $\tau$, and compare with the baselines. Defaults of NN, $B$, $\alpha$, $\tau$ are 3535, 40004000, $0.3$ and 66.
@@
To illustrate how the TTSnap parameters contribute to the reward of the final selected image, we provide a one-stage pruning analysis in Fig. 5 with PickScore. From Fig. 4(a), increasing the number of candidates (NN) consistently improves the final rewards, and our finetuned noise-aware reward model converges faster as NN grows compared to baseline models, since it provides more accurate intermediate reward estimates. Fig. 4(b) highlights another key reason why TTSnap outperforms best-of-NN. With pruning, we can explore many more candidates under the same compute budget. In Fig. 4(c) and Fig. 4(d), we examine the effect of the retention ratio $\alpha$ and the pruning timestep $\tau$. A larger $\alpha$ and a later pruning stage $\tau$ both lead to better results, since they reduce the risk of discarding promising trajectories. As indicated by our cost formulation in Eq. 7, increasing $\alpha$ or $\tau$ raises the computational cost linearly, while the performance improves quickly at first but then gradually flattens out, resulting in diminishing marginal gains.
@@
TTSnap performance can be further improved by extending the algorithm to a multi-stage pruning setup. We search for the optimal parameter settings across different numbers of pruning stages, as shown in Tab. 1, and performance is evaluated using Eq. 11 within our budget range. According to Eq. 7 and Eq. 8, adding more stages saves budget and allows a broader exploration of candidates, but it also accumulates more reward estimation errors. As a result, when using more stages, the retention ratios $\alpha$ should be set slightly higher, especially in later timesteps when fewer candidates remain. While increasing a certain number of pruning stages improves performance, too many stages become counterproductive, since the verification cost would be highly increased, eventually reducing performance under budget forcing. For comparison, we also search for the best parameter settings for rewards without noise-aware finetuning (NARF) and plot them with the best-of-NN baseline in Fig. 5(a). As shown in the figure, while best-of-NN requires 4000 TFLOPs to reach that reward level, our method achieves it with only 2400 TFLOPs.
Table 1: Optimal search parameters for multi-stage pruning with our noise-aware PickScore reward model. We report the relative performance improvement across a range of budgets $\omega$, with respect to the best-of-NN strategy.
@@
Table 2: Comparison of the best-of-NN (BoN) strategy, pruning without NARF (TTSp), and pruning with NARF (TTSnap) under single-reward and combined-reward settings. The top row indicates the reward model used by the algorithm and the second row indicates the reward model used for evaluation. We report the final reward values for compute budgets $B\!=\!1\mathrm{k},2\mathrm{k},4\mathrm{k}$ TFLOPs, along with the relative gain $\omega$ over best-of-NN. $r(B\!\rightarrow\!0)$ denotes the expected reward of sampling one image per prompt at random.
@@ 6.3 Discussion
Given the above observation, more global exploration generally leads to stronger TTS performance. Our method focuses on maximizing the number of explored candidates within a given budget. In contrast, some approaches consider an orthogonal direction that prioritizes local refinement around promising candidates [tts_ma, tts_greedy, dynamic_search]. Their motivation is that global search over the initial noise samples, as used in best-of-NN and our method, carries the risk of reward hacking [reward_hack]. However, this risk can be effectively controlled by limiting the search size NN or by using multiple reward models to provide verification signals from different perspectives. As demonstrated in Tab. 2, reducing NN or combining multiple rewards helps prevent the output images from overfitting to any single reward model.
@@ A.1 TTSnap Performance on SDXL
Table 4: Evaluation of TTSnap performance on SDXL. As for Flux, we compare the best-of-NN (BoN) strategy, pruning without NARF (TTSp), and pruning with NARF (TTSnap) under single-reward and combined-reward settings. The top row indicates the reward model used by the algorithm and the second row indicates the reward model used for evaluation. We report the final reward values for compute budgets $B\!=\!1\mathrm{k},2\mathrm{k},4\mathrm{k}$ TFLOPs, along with the relative gain $\omega$ over best-of-NN. $r(B\!\rightarrow\!0)$ denotes the expected reward of sampling one image per prompt at random.
@@ Appendix D Further Details
where $C$ and $D$ denote the numbers of concordant and discordant pairs, respectively. A pair $(i,j)$ is concordant if the two rankings agree on their relative order and discordant otherwise. The number of discordant pairs corresponds to the number of swaps that bubble sort would perform to transform one ranking into the other. Since Kendall’s $\tau$ depends solely on the ordering rather than the magnitude of the rank values, it penalizes only ordinal inconsistencies. Identical rankings receive a score of 11, completely reversed rankings receive 1-1, and independent random rankings have an expected value of 0.

### v2

@@ 1 Introduction
Figure 1: Comparison of the best-of-$N$ TTS algorithm with our TTSnap algorithm under an equivalent compute budget. By applying early-stage pruning, TTSnap can explore a much larger pool of candidates (shown as polygons). Candidate colors indicate reward values from low (light yellow) to high (dark red). Each candidate is denoised from left to right under different initial noise. The magnifying glass represents the reward models, where the noise-aware variants are finetuned using our self-distillation strategy.
@@ 3 Preliminaries
and may not be differentiable with respect to the inference trajectory. The search algorithm aims to identify the most promising candidates that maximize the verifier’s reward scores. The algorithm $f$ takes the verifier $\psi\in\Psi$, the diffusion model $\phi\in\Phi$, and $N$ candidate image samples with text conditions as input and returns an image:
@@
Increasing the number of candidates $N$ increases the probability of finding a high-reward sample, but also increases the inference-time compute cost. Among various search algorithms, best-of-$N$ is the most straightforward yet effective one [sana1_5, tts_ma]. This algorithm independently samples $N$ Gaussian noise, denoises each, and verifies their rewards. Its core advantage lies in enabling global exploration across the initial noise space. Other strategies, such as greedy neighborhood search [tts_ma, tts_greedy] or trajectory-level search [fk, tts_ma, dynamic_search, beam_vseq, tts_kim], primarily focus on finding locally optimal candidates. In this paper, we mainly focus on improving TTS over best-of-$N$, since a more global search, as discussed in Sec. 6.3, typically yields better performance. While some methods scale performance by refining the conditioning signals [reflect_perfect, reflectdit, tts_kim_p], we focus our study to the setting where the text condition remains fixed.
@@ 4 The TTSnap Algorithm
The primary limitation of the best-of-$N$ strategy lies in its high verification cost, as each candidate must be fully denoised before reward evaluation. This constraint severely restricts the breadth of candidate exploration under a fixed computational budget. To address this inefficiency, we propose a pruning strategy, outlined in Algorithm 1, that verifies and eliminates unpromising candidates early in the denoising process. Since the reward models operate in the image domain, we approximate the image corresponding to an intermediate latent by estimating its posterior mean, detailed in Sec. 3, and decoding it via the variational auto-encoder (VAE) decoder. The intermediate images are then scored by the reward model, and candidates with low rewards are pruned according to a predefined retention ratio. This process is iterated until the final step, where the remaining candidates are fully denoised, and the one yielding the highest reward is selected as the output. The set of verification timesteps is denoted as $\mathcal{T}=\{\tau_{1},\dots,\tau_{m-1}\}\subseteq[1,M-1]$, where $\tau_{m-1}=M-1$. $M$ is the final timestep. The set of corresponding retention ratios is $\mathcal{A}=\{\alpha_{1},\dots,\alpha_{m-1}\}$, where $0<\alpha_{i}<1$ is the fraction of candidates that will be retained at timestep $\tau_{i}$.
@@ 12    $\mathcal{X}\leftarrow\{\mathcal{X}_{i}\mid i\in\operatorname*{arg\,topk}(\mathcal{R},k)\}$
By avoiding the full denoising of every candidate, such an algorithm can substantially reduce the overall compute cost. However, it introduces additional verification overhead: VAE decoding and reward computation. Let the computational cost of a single denoising step for one candidate be $B_{d}$, and that of a single verification be $B_{v}$. For $N$ initial candidates, the total compute budget $B$ can be expressed as
@@
Under the same compute budget $B$, our $N$ is much larger than that of the original best-of-$N$ strategy: $N=\lfloor B/(MB_{d}+B_{v})\rfloor$.
@@
The primary goal of TTS is to maximize the generated image quality with minimal computational cost. Existing TTS methods are typically evaluated under a fixed computation budget or by visualizing the relationship between the final reward and computational cost [dynamic_search, tts_ma, reflectdit]. Such evaluations are often indirect. Hence, we introduce a new metric that quantifies performance improvement as the computation budget increases. Let $f(N)$ denote the final image obtained from algorithm $f$ with $N$ noise candidates, as described in Eq. 6, and let $\psi(f(N))$ represent its corresponding reward score. The number of explored candidates $N$ can be determined from a given computation budget $B$, such as Eq. 8 for our method. Over a budget range $B^{-},B^{+}]$, we define the incremental performance gain with respect to budget $B$ as
@@ Results.
We construct a validation set using 200 prompts unseen during training, generating 80 samples per prompt along with all corresponding intermediate images. For the purpose of TTS, our goal is to ensure that reward predictions at early inference stages preserve a ranking similar to the final clean-image counterparts. To measure this rank consistency, we compute Kendall’s $\tau$ coefficient [kendall] between the noise-aware reward scores at early timesteps and the final-step reward scores for each prompt, and report the average correlation across all prompts on the validation set. Kendall’s $\tau$ measures the ordinal agreement between two ranked lists, taking values from $1$ (identical order) to 1-1 (complete reversal). It reflects the normalized number of pairwise swaps needed to turn one ranking into the other, independent of the value magnitudes. As shown in Fig. 3, our noise-aware reward self-distillation consistently improves early-stage ranking agreement across all three reward models compared with their un-finetuned baselines. This substantial gain in rank consistency makes our early-step pruning framework more reliable and effective. A further analysis of NARF and a comparison with its alternatives are presented in Appendix Appendix B.
@@ 6.2 TTSnap Experiments
We take the fine-tuned noise-aware reward model from the previous section and evaluate our TTSnap algorithm on the validation prompts. The computation budget is measured in TFLOPs. We cap the maximum budget at 4,000 TFLOPs, which approximately equals 400 forward passes of the Flux diffusion transformer. Our main evaluation metric is $w$ as defined in Eq. 11, representing the increase ratio of the integral of reward improvement over budget, with reference to the best-of-$N$ baseline. To mitigate randomness and ensure a fair comparison across different TTSnap configurations, we generate a pool of over 200 candidate trajectories for each prompt and constrain all search procedures to operate within this shared pool. For each search, we iterate 100 times over different seeds and report the average reward of the final selected image.
@@
Figure 4: Analysis of the effect of different variables on the PickScore reward or number of explored samples NN, with respect to (a) NN, (b) the compute budget $B$, (c) the retention ratio $\alpha$, and (d) the pruning timestep $\tau$, and compare with the baselines. Defaults of $N$, $B$, $\alpha$, $\tau$ are 3535, 40004000, $0.3$ and $6$.
@@
To illustrate how the TTSnap parameters contribute to the reward of the final selected image, we provide a one-stage pruning analysis in Fig. 5 with PickScore. From Fig. 4(a), increasing the number of candidates ($N$) consistently improves the final rewards, and our finetuned noise-aware reward model converges faster as $N$ grows compared to baseline models, since it provides more accurate intermediate reward estimates. Fig. 4(b) highlights another key reason why TTSnap outperforms best-of-$N$. With pruning, we can explore many more candidates under the same compute budget. In Fig. 4(c) and Fig. 4(d), we examine the effect of the retention ratio $\alpha$ and the pruning timestep $\tau$. A larger $\alpha$ and a later pruning stage $\tau$ both lead to better results, since they reduce the risk of discarding promising trajectories. As indicated by our cost formulation in Eq. 7, increasing $\alpha$ or $\tau$ raises the computational cost linearly, while the performance improves quickly at first but then gradually flattens out, resulting in diminishing marginal gains.
@@
TTSnap performance can be further improved by extending the algorithm to a multi-stage pruning setup. We search for the optimal parameter settings across different numbers of pruning stages, as shown in Tab. 1, and performance is evaluated using Eq. 11 within our budget range. According to Eq. 7 and Eq. 8, adding more stages saves budget and allows a broader exploration of candidates, but it also accumulates more reward estimation errors. As a result, when using more stages, the retention ratios $\alpha$ should be set slightly higher, especially in later timesteps when fewer candidates remain. While increasing a certain number of pruning stages improves performance, too many stages become counterproductive, since the verification cost would be highly increased, eventually reducing performance under budget forcing. For comparison, we also search for the best parameter settings for rewards without noise-aware finetuning (NARF) and plot them with the best-of-$N$ baseline in Fig. 5(a). As shown in the figure, while best-of-$N$ requires 4000 TFLOPs to reach that reward level, our method achieves it with only 2400 TFLOPs.
Table 1: Optimal search parameters for multi-stage pruning with our noise-aware PickScore reward model. We report the relative performance improvement across a range of budgets $\omega$, with respect to the best-of-$N$ strategy.
@@
Table 2: Comparison of the best-of-$N$ (BoN) strategy, pruning without NARF (TTSp), and pruning with NARF (TTSnap) under single-reward and combined-reward settings. The top row indicates the reward model used by the algorithm and the second row indicates the reward model used for evaluation. We report the final reward values for compute budgets $B\!=\!1\mathrm{k},2\mathrm{k},4\mathrm{k}$ TFLOPs, along with the relative gain $\omega$ over best-of-NN. $r(B\!\rightarrow\!0)$ denotes the expected reward of sampling one image per prompt at random.
@@ 6.3 Discussion
Given the above observation, more global exploration generally leads to stronger TTS performance. Our method focuses on maximizing the number of explored candidates within a given budget. In contrast, some approaches consider an orthogonal direction that prioritizes local refinement around promising candidates [tts_ma, tts_greedy, dynamic_search]. Their motivation is that global search over the initial noise samples, as used in best-of-$N$ and our method, carries the risk of reward hacking [reward_hack]. However, this risk can be effectively controlled by limiting the search size $N$ or by using multiple reward models to provide verification signals from different perspectives. As demonstrated in Tab. 2, reducing $N$ or combining multiple rewards helps prevent the output images from overfitting to any single reward model.
@@ A.1 TTSnap Performance on SDXL
Table 4: Evaluation of TTSnap performance on SDXL. As for Flux, we compare the best-of-$N$ (BoN) strategy, pruning without NARF (TTSp), and pruning with NARF (TTSnap) under single-reward and combined-reward settings. The top row indicates the reward model used by the algorithm and the second row indicates the reward model used for evaluation. We report the final reward values for compute budgets $B\!=\!1\mathrm{k},2\mathrm{k},4\mathrm{k}$ TFLOPs, along with the relative gain $\omega$ over best-of-NN. $r(B\!\rightarrow\!0)$ denotes the expected reward of sampling one image per prompt at random.
@@ Appendix D Further Details
where $C$ and $D$ denote the numbers of concordant and discordant pairs, respectively. A pair $(i,j)$ is concordant if the two rankings agree on their relative order and discordant otherwise. The number of discordant pairs corresponds to the number of swaps that bubble sort would perform to transform one ranking into the other. Since Kendall’s $\tau$ depends solely on the ordering rather than the magnitude of the rank values, it penalizes only ordinal inconsistencies. Identical rankings receive a score of $1$, completely reversed rankings receive 1-1, and independent random rankings have an expected value of 0.